# Low-Energy On-Device Personalization for MCUs

### v3

@@ 1. Introduction
Microcontroller Units (MCUs) are cost-effective and energy-efficient, enabling wide opportunities for Tiny Machine Learning (TinyML) (dutta2021tinyml, ). For example, MCUs can be used in wearable devices for health monitoring (haghi2017wearable, ), smart home appliances for energy management (abadade2023comprehensive, ), and industrial sensors for predictive maintenance (kanawaday2017machine, ). However, the diverse range of deployment scenarios for MCUs makes them vulnerable to data drift issues.
Data drift (liang2023comprehensive, ), when the distributions of source and target data diverge, can result in a drop in model performance. To solve this challenge, common personalization approaches such as multi-task learning (zhang2018overview, ; hu2021unit, ) and zero-shot learning (pourpanah2022review, ), as well cloud-edge methods like federated learning (mcmahan2017communication, ; hao2020time, ; chai2021fedat, ), have been proposed. Nevertheless, these approaches have their limitations: multi-task learning and zero-shot learning can struggle with the complexity and variability of local data, while cloud-edge strategies rely on frequent and stable communication between the cloud and local resources, often requiring significant computational resources. Therefore, personalizing models locally is essential to maintain their effectiveness.
On-device personalization, performed entirely by the local device without cloud communication, improves efficiency and reduces latency (goldenberg2021personalization, ). $TinyTL$ (cai2020tinytl, ) and $SparseUpdate$ (lin2022device, ) both achieve personalization on MCUs and have been deployed and tested with relatively advanced models. However, our experiments found that despite making significant contributions to model performance, they consume significant energy (detailed in Section 6). $TinyTL$ adds lite-residual modules to the model backbone and only trains these added modules on the MCU. Although these modules have been carefully designed, they still require a significant amount of energy for training. $SparseUpdate$ only trains the selected layer/channel to improve the efficiency. However, the layer/channel-searching requires access to an entire local dataset, which is not applicable in real scenarios. It also requires thousands of computationally intensive search (lin2022device, ) or pruning (profentzas2022minilearn, ) processes, resulting in long training times of up to ten days (kwon2023tinytrain, ). Moreover, these methods focus solely on local training and overlook local inference, which impacts the energy consumption of the MCU during long-term operation.
@@
Contribution. In this study, we introduce $MicroT$, a low-energy model personalization framework for ultra-resource-constrained MCUs  [^1] We will release our code and design as open source upon acceptance.. The goal of $MicroT$ is to sustain personalization performance while reducing its energy consumption. We consider that model personalization includes both training and inference and therefore explored model performance and energy consumption in both of these stages. Our contributions are summarized as follows:
@@ 3.1. Overview
MicroT splits the model into a feature extractor and classifier. $MicroT$ adopts the $TTL$ approach by dividing the model into a feature extractor and a classifier, guided by energy and computational requirements. We found that the performance improvements obtained by existing TTL approaches for MCUs, such as $TinyTL$ and $SparseUpdate$, are at the cost of increased energy and computational demands (detailed in Section. 6), undermining the advantages of MCUs. While the ‘feature extractor + classifier’ paradigm results in inferior model performance, it offers the lowest energy consumption and computational demands. We believe that, from a practical standpoint, the most direct approach for optimization is to retain this paradigm and optimize its performance.
@@
Given the limited memory of MCUs, we can only deploy lightweight models, such as MCUNet (lin2020mcunet, ). In our experiments, we found that directly applying $SSL$ to these lightweight models resulted in limited performance improvements, with the model capacity causing a bottleneck (azizi2021big, ) (detailed in Section 7). Therefore, to further optimize the feature extractor, we combined SSL and KD (SSKD), as shown in Fig. 3.1 Stage 1. Specifically, we first use SSL to train a large model on a public dataset, then use the large model to obtain a generative dataset, and finally, train a smaller feature extractor on both the generative dataset and the public dataset.
@@ 3.2. Self-Supervised Knowledge Distillation
where $g_{i}$ and $e_{i}$ represent the output logits of the guide model and the exploration model, respectively, and $\tau_{g}$ and $\tau_{e}$ are temperature parameters that adjust the sensitivity of the $softmax$. This setup encourages the exploration model to learn the behavior of the guide model by processing different data views, while the guide model’s EMA strategy ensures it gradually adapts and improves on its ability to capture data representations.
@@ 3.3. Model Segmentation and Joint Training
where $F_{\text{score}}$ is the score used to evaluate the model segmentation point. $R_{A}^{\text{norm}}$ represents the normalized ratio of accuracy loss, which measures the decrease in the part model’s accuracy relative to the full model’s accuracy. $R_{M}^{\text{norm}}$ is the normalized Multiply–Accumulate ($MAC$) reduction ratio, which quantifies the decrease in the number of $MACs$ in the part model’s relative to the number of $MACs$ in the full model. $G$ is the normalized gain ratio, representing the trade-off between the increased accuracy and $MACs$ from the part model split.
@@ 4.4. Baselines
For model performance, we compare $MicroT$ with four baselines: (1) $None$ does not perform any on-device training; (2) $TTL$ is traditional transfer learning, which only trains the task head (wu2020emo, ); (3) $TinyTL$ adds lite-residual modules to the model backbone and freezes the backbone during training (cai2020tinytl, ); (4) $SparseUpdate$ statically selects the layers and channels to update by multiple gradient computation, and then trains these layers/channels online (lin2022device, ). For a fair comparison, we use the same number of training epochs for all methods and grid search to determine the optimal learning rate. To test the optimal performance of the methods, we utilize the complete dataset and split it into training/testing/validation dataset with a ratio of 8:1:1. Although $MicroT$’s stage-decision ratio is configurable, for easier comparison with the baseline(s), we utilize the standard stage-decision ratio of 0.5.
@@ 5.2. How Does the Performance of MicroT Compare to the Baselines?
We set $MicroT$ to the standard stage-decision rate of 0.5 in order for easier compared to our the baselines, the results are shown in Table. 3. It can be observed that $None$ has the lowest accuracy across all experiments. This indicates that when deploying models locally on MCUs, personalization is necessary. Compared to $None$, $TTL$ shows a small improvement in accuracy, which suggests that merely training the classifier alone is insufficient to fully adapt towards the local personalized data. However, we do not abandon the ’train only the classifier’ paradigm as it offers the lowest computational requirements and energy consumption for MCUs (detailed in Section. 6). Instead, to better utilize the cloud resource available, we look towards improving the quality of the feature extractor.
@@ 5.3. How Many Samples Do We Need to Determine the Confidence Thresholds?
The confidence threshold dictates the stage-decision ratio, determining the proportion of samples processed only by the part model. The threshold setting is based on the confidence score sequence from the part model. Specifically, the trained part model processes $N$ samples and generates a confidence score sequence. We then select the median of this sequence to obtain the standard stage-decision ratio of 0.5. Ideally, selecting the median of the entire test samples’ confidence sequence sets the stage-decision ratio to 0.5, meaning 50% of samples are processed only by the part model. However, in real-world applications, it is unreasonable to waste too much time finding this threshold. Therefore, to understand if it is efficient to determine the threshold locally, we process various quantities of samples to calculate their medians, and compare these medians’ corresponding stage-decision ratios and model performance. We do this to understand the number of local samples needed to accurately determine the median and standard stage-decision ratio. The results are shown in Table.  4.
@@ 6. System Cost
$MicroT$ supports two phases on the MCU: classifier training and inference. For the training energy cost, we compare $MicroT$ with $TTL$ and $TinyTL$. We do not compare $SparseUpdate$ as it takes 1̃0 days to search for the layers and channels to update and train the model, and this is inhibitively energy-consuming. For the energy cost of inference, we select $TTL$ as our baseline ($i.e.$, full model inference), representing the inference approach used in most on-device personalization studies. We select the Plant (goeau2017plant, ) and Bird datasets (WahCUB_200_2011, ), as these datasets contain the smallest and largest number of classes respectively. To efficiently analyze the energy cost, we use the average of these two datasets.
@@
We utilize the Monsoon Power Monitor (msoon, ) to measure the time and average power of training, then calculate their energy consumption (as described in Section. 4.3). When measuring the energy cost of local training, we train the classifier on 100 local data samples for 40 epochs ($i.e.$, 4,000 iterations total). We ignore the energy consumption of other MCU components or operations and only consider the energy consumption of local training, including data preprocessing, feature extractor inference, and classifier training. The results are shown in Fig.  7.
@@
Table 6. The energy costs ($kJ$) of the MCU local inference with various stage-decision ratio.
@@ 8. Discussion
Limitations. We have currently tested the performance of $MicroT$ on a limited number of models, but its versatility enables it to be easily applied to other lightweight models, such as SqueezeNet (iandola2016squeezenet, ), MobileNet (howard2017mobilenets, ), and models generated by Neural Architecture Search (NAS). $MicroT$ can also be easily combined with other model compression methods, such as quantization (xiao2023smoothquant, ; jacob2018quantization, ) and pruning (zhu2017prune, ; liu2020pruning, ).
@@
Unsupervised Personalization. $MicroT$ relies on labeled data for local training, similar to most on-device personalization approaches (cai2020tinytl, ; lin2022device, ; kwon2023tinytrain, ). However, exploring unsupervised training, such as distance-based k-means classifiers (wu2020emo, ), is an interesting direction. We conducted experiments with traditional k-means classifiers and found they performed well on simple datasets. However, the traditional k-means classifiers did not perform well on complex datasets, likely due to the large number of categories and dense feature space. Nevertheless, improving the differentiation of categories in the feature space through methods such as contrastive learning (chen2020simple, ) may enhance the performance of unsupervised personalization. Such advancements would broaden the applicability of $MicroT$.
@@
Multi-modal Feature. We only evaluated $MicroT$ using image data as an example. However, the ability to manage diverse signal sources, like environmental sensors (daghero2022human, ; kaewmard2014sensor, ), is also essential for MCUs. Recent studies have investigated data and feature fusion techniques for multi-modal data analysis on edge devices (wang2022am3net, ; ouyang2022cosmo, ). $MicroT$ could integrate its feature extractor with these fusion methods to map multi-modal data into a unified feature space. This expansion not only widens $MicroT$’s utility but also enhances its robustness. Other tasks, such as person detection, vehicle counting, and audio classification, are also interesting future research directions and applications.

### v4

@@ 1. INTRODUCTION
Microcontroller Units (MCUs) are cost-effective and energy-efficient, enabling wide opportunities for Tiny Machine Learning (TinyML) (dutta2021tinyml, ). For example, MCUs can be used in wearables for health monitoring (haghi2017wearable, ), smart home for energy management (abadade2023comprehensive, ), and industrial sensors for predictive maintenance (kanawaday2017machine, ). However, the diverse range of deployment scenarios for MCUs makes them vulnerable to data drift issues.
Data drift (liang2023comprehensive, ), when the distributions of source and target data diverge, can result in a drop in model performance. To solve this challenge, common personalization approaches such as multi-task learning (zhang2018overview, ; hu2021unit, ), zero-shot learning (pourpanah2022review, ), and cloud-edge methods like federated learning (mcmahan2017communication, ; hao2020time, ; chai2021fedat, ), have been proposed. Nevertheless, these approaches have limitations: multi-task learning and zero-shot learning can struggle with the complexity and variability of local data, while cloud-edge strategies rely on frequent and stable communication between the cloud and local resources, often requiring significant computational resources. Therefore, personalizing models locally is essential to maintain their effectiveness.
On-device personalization, performed entirely by the local device without cloud communication, improves efficiency and reduces latency (goldenberg2021personalization, ). $TinyTL$ (cai2020tinytl, ) and $SparseUpdate$ (lin2022device, ) both achieve personalization on MCUs and have been deployed and tested with advanced models. However, our experiments found that despite making significant contributions to model performance, they consume significant energy (detailed in Section 6). $TinyTL$ adds lite-residual modules to the backbone and only trains these added modules on the MCU. Although these modules have been carefully designed, they still require a significant amount of energy for training. $SparseUpdate$ only trains the selected layer/channel to improve the efficiency. However, the layer/channel-searching requires access to an entire local dataset, which is not applicable in real scenarios. It also requires thousands of computationally intensive search (lin2022device, ) or pruning (profentzas2022minilearn, ) processes, resulting in long training times of up to ten days (kwon2023tinytrain, ). Moreover, these methods focus solely on local training and overlook local inference, which impacts the energy consumption of the MCU during long-term operation.
@@
Contribution. In this study, we introduce $MicroT$, a low-energy model personalization framework for ultra-resource-constrained MCUs  [^1] https://github.com/yushan-huang/MicroT. The goal of $MicroT$ is to sustain personalization performance while reducing its energy consumption. We consider that model personalization includes both training and inference and therefore explored model performance and energy consumption in both of these stages. Our contributions are summarized as follows:
@@ 3.1. Overview
Feature extractor + classifier paradigm. $MicroT$ adopts the $TTL$ approach, dividing the model into a feature extractor and a classifier, guided by energy and computational requirements. We found that the performance improvements obtained by existing MCU $TTL$ approaches, such as $TinyTL$ and $SparseUpdate$, are at the cost of increased energy and computational demands (detailed in Section. 6), undermining the advantages of MCUs. While the ‘feature extractor + classifier’ paradigm results in inferior model performance, it offers the lowest energy consumption and computational demands. We believe that, from a practical standpoint, the most direct approach for optimization is to retain this paradigm and optimize its performance.
@@
Given the limited memory of MCUs, we can only deploy lightweight models, such as MCUNet (lin2020mcunet, ). In our experiments, we found that directly applying $SSL$ to these lightweight models resulted in limited performance improvements, with the model capacity causing a bottleneck (azizi2021big, ) (detailed in Section 7). Therefore, to optimize the feature extractor, we combined SSL and KD (SSKD), as shown in Fig. 3.1 Stage 1. Specifically, we first use SSL to train a large model on a public dataset, then use the large model to obtain a generative dataset, and finally, train a smaller feature extractor on the generative dataset and the public dataset.
@@ 3.2. Self-Supervised Knowledge Distillation
where $g_{i}$ and $e_{i}$ represent the output logits of the guide model and the exploration model. $\tau_{g}$ and $\tau_{e}$ are temperature parameters that adjust the sensitivity of the $softmax$. This setup encourages the exploration model to learn the behavior of the guide model by processing different data views, while the guide model’s EMA strategy ensures it gradually adapts and improves the data representation capturing ability.
@@ 3.3. Model Segmentation and Joint Training
where $F_{\text{score}}$ is the score used to evaluate the segmentation point. $R_{A}^{\text{norm}}$ represents the normalized ratio of accuracy loss, which measures the decrease in the part model’s accuracy relative to the full model’s accuracy. $R_{M}^{\text{norm}}$ is the normalized Multiply–Accumulate ($MAC$) reduction ratio, which quantifies the decrease in the number of $MACs$ in the part model’s relative to the number of $MACs$ in the full model. $G$ is the normalized gain ratio, representing the trade-off between the increased accuracy and $MACs$ from the part model.
@@ 4.4. Baselines
For model performance, we compare $MicroT$ with four baselines: (1) $None$ does not perform any on-device training, and the local classifier is randomly initialized; (2) $TTL$ is traditional transfer learning, which only trains the task head (wu2020emo, ); (3) $TinyTL$ adds lite-residual modules to the model backbone and freezes the backbone during training (cai2020tinytl, ); (4) $SparseUpdate$ statically selects the layers and channels to update by multiple gradient computation, and then trains these layers/channels online (lin2022device, ). To demonstrate the improvements brought by $MicroT$, we use the same feature extractor (MCUNet and ProxylessNAS) for $MicroT$ and all baselines. We also use the same number of training epochs for all methods and grid search to determine the optimal learning rate. To test the optimal performance of the methods, we utilized the complete dataset and split it into training/testing/validation dataset with a ratio of 8:1:1. Although $MicroT$’s stage-decision ratio is configurable, for easier comparison with the baseline(s), we utilize the standard stage-decision ratio of 0.5.
@@ 5.2. How Does the Performance of MicroT Compare to the Baselines?
We set $MicroT$ to the standard stage-decision rate of 0.5 in order for easier compared to our the baselines, the results are shown in Table. 3. It can be observed that $None$ has the lowest accuracy across all experiments. This indicates that when deploying models locally on MCUs, personalization is necessary. Compared to $None$, $TTL$ shows significant improvement. This indicates that the transfer learning paradigm is an effective method for solving local personalization problems. Compared to $TinyTL$ and $SparseUpdate$, $TTL$ still leaves room for improvement, which suggests that merely training the classifier is insufficient to fully adapt towards the local personalized data. However, we do not abandon the ’training only the classifier’ paradigm, as it offers the lowest computational requirements and energy consumption for MCUs (detailed in Section. 6). Instead, to better utilize the cloud resource available, we look towards improving the quality of the feature extractor.
@@ 5.3. How Many Samples Do We Need to Determine the Confidence Thresholds?
The confidence threshold dictates the stage-decision ratio, determining the proportion of samples processed only by the part model. The threshold setting is based on the confidence score sequence from the part model. Specifically, the trained part model processes $N$ samples and generates a confidence score sequence. The median of this sequence is then selected to obtain the standard stage-decision ratio of 0.5. Ideally, selecting the median of the entire test samples’ confidence sequence sets the stage-decision ratio to 0.5, meaning 50% of samples are processed only by the part model. However, in real-world applications, it is unreasonable to waste too much time finding this threshold. Therefore, to understand if it is efficient to determine the threshold locally, we randomly select various quantities of samples (regardless of category) from the entire dataset to calculate medians, and compare these medians’ corresponding stage-decision ratios and model performance. We do this to understand the number of local samples needed to accurately determine the median and standard stage-decision ratio. The results are shown in Table.  4.
@@ 6. SYSTEM COST
$MicroT$ has two phases on the MCU: classifier training and inference. For the training energy cost, we compare $MicroT$ with $TTL$ and $TinyTL$. We do not compare $SparseUpdate$ as it takes 1̃0 days to search for the layers and channels to update and train the model, and this is inhibitively energy-consuming. For the energy cost of inference, we select $TTL$ as our baseline ($i.e.$, full model inference), representing the inference approach used in most on-device personalization studies. We select the Plant (goeau2017plant, ) and Bird datasets (WahCUB_200_2011, ), as these datasets contain the smallest and largest number of classes respectively. To efficiently analyze the energy cost, we use the average of these two datasets.
@@
We utilize the Monsoon Power Monitor (msoon, ) to measure the time and average power of training, then calculate their energy consumption (as described in Section. 4.3). To facilitate comparison and measurement of the energy consumption, we use the example of training 100 local data samples for 40 epochs ($i.e.$, 4,000 iterations total). We ignore the energy consumption of other MCU components or operations and only consider the energy consumption of local training, including data preprocessing, feature extractor inference, and classifier training. The results are shown in Fig.  7.
@@
Table 6. The model performances (%) with various stage-decision ratios.
Model Method Pet Bird Plant Avg.
MCUNet MicroT_0.25 84.34 56.05 53.87 64.75
MicroT_0.75 76.68 47.87 51.65 58.73
ProxylessNAS MicroT_0.25 82.35 51.97 54.04 62.79
MicroT_0.75 76.44 46.59 50.36 57.80
Table 7. The energy costs ($kJ$) of the MCU local inference with various stage-decision ratios.
@@ 7. ABLATION STUDY
Model performance with various stage-decision ratio. In Section. 5, we have introduced the model performance of $MicroT$ with ratio of 0.5. To provide a clear demonstration of the performance trend, we further show the model performance with ratios of 0.25 and 0.75. The results are shown in Table. 6.
@@ 8. DISCUSSION
In this section, we discuss some analysis and limitations of the methodology and experiments in this study, as well as the future works.
(1) Application scenarios. In this study, we primarily conducted experiments in the image domain, which offers many open-source models/datasets, making it convenient for us to validate $MicroT$. However, $MicroT$ has several application scenarios in energy-sensitive environments. For instance, battery-free underwater devices powered by acoustic energy that use low-power cameras to detect and identify marine species (afzal2022battery, ); solar-powered low-power sensor networks for monitoring wildlife (hart2020precision, ); environmental monitoring devices powered by wind or water energy that require long-term self-sustainability (samijayani2023solar, ); and implantable medical devices powered by bioenergy (jiang2020emerging, ). These practical applications can all leverage the technologies related to $MicroT$.
@@
(2) The experimental models and MCUs. We have currently tested the performance of $MicroT$ on a limited number of models, but its versatility enables it to be easily applied to other lightweight models, such as SqueezeNet (iandola2016squeezenet, ), MobileNet (howard2017mobilenets, ), and models generated by Neural Architecture Search (NAS). $MicroT$ can also be easily combined with other model compression methods, such as quantization (xiao2023smoothquant, ; jacob2018quantization, ) and pruning (zhu2017prune, ; liu2020pruning, ).
(3) Stability of model segmentation. One of the objectives of the feature extractor is to enhance its generalization, which means that the feature extractor can also exhibit high generalization on local datasets. Therefore, the global optimal segmentation point is potentially to be the optimal or near-optimal point on the local dataset as well. However, if the local optimal point differs from the global point, our strategy is to conservatively adopt the global point because it is a more stable and general choice. Even if the optimal point varies slightly on the local dataset, the overall performance could still meet requirements.
(4) The ‘feature extractor + classifier’ paradigm. This paradigm has advantages in energy efficiency, which is the core focus of $MicroT$ for ultra-low-power devices. However, to reduce resource consumption, local training only involves the classifier, which places high demands on the generalization of the feature extractor. Nevertheless, this is a necessary design consideration on performance and energy efficiency. $MicroT$ remains effective in most real-world application scenarios because the generalization of the feature extractor can be ensured: first, the rapid development of smart devices has provided a wealth of publicly available training datasets; second, we can leverage technologies such as SSL.
(5) Re-training classifiers. Ideally, once the classifier is trained locally and meets task requirements, it can stop training and be used for a long time. However, in practice, local data often undergoes some degree of drift, such as changes in lighting conditions due to seasonal variations. While the general feature extractor can handle minor local drifts, in some extreme cases, further fine-tuning of the classifier may still be necessary. We believe that this fine-tuning is usually quick and straightforward, as local data distribution changes are often gradual rather than sudden.
@@
(6) Unsupervised Personalization. $MicroT$ relies on labeled data for local training, similar to most on-device personalization approaches (cai2020tinytl, ; lin2022device, ; kwon2023tinytrain, ). However, exploring unsupervised training, such as distance-based k-means classifiers (wu2020emo, ), is an interesting direction. We conducted experiments with traditional k-means classifiers and found they performed well on simple datasets. However, the traditional k-means classifiers did not perform well on complex datasets, likely due to the large number of categories and dense feature space. Nevertheless, improving the differentiation of categories in the feature space through methods such as contrastive learning (chen2020simple, ) may enhance the performance of unsupervised personalization. Such advancements would broaden the applicability of $MicroT$.
@@
(7) Multi-modal Feature. We only evaluated $MicroT$ using image data as an example. However, the ability to manage diverse signal sources, like environmental sensors (daghero2022human, ; kaewmard2014sensor, ), is also essential for MCUs. Recent studies have investigated data and feature fusion techniques for multi-modal data analysis on edge devices (wang2022am3net, ; ouyang2022cosmo, ). $MicroT$ could integrate its feature extractor with these fusion methods to map multi-modal data into a unified feature space. This expansion not only widens $MicroT$’s utility but also enhances its robustness. Other tasks, such as person detection, vehicle counting, and audio classification, are also interesting future research directions and applications.